OpenStudy (anonymous):

what is the exact value of 7q^4 / sqrt q^15

OpenStudy (aravindg):

first take q out of square root in denominator

OpenStudy (anonymous):

\[\frac{ 7q^4 }{\sqrt{q^{15}} } = 7q^{4 - (15/2)} = 7q^{-7/2} =\frac{ 7 }{ q^{7/2} }\]

OpenStudy (anonymous):

but these are the only options. lol

OpenStudy (anonymous):

\[\frac{ 7 }{ q^{7/2} } \times\frac{ \sqrt{q} }{ \sqrt{q} } = \frac{7\sqrt{q} }{\sqrt{q^{7}}\sqrt{q}} = \frac{ 7\sqrt{q} }{ q^{8/2} } = \frac{7\sqrt{q}}{q^4}\]